OpenStudy (anonymous):

What is the 35th term of the arithmetic sequence where a1 = 13 and a17 = -67?

OpenStudy (nurali):

a(n) = a(1) + (n - 1)d a(17) = -67 a(1) + (17 - 1)d = -67 13 + 16d = -67 16d = -67 - 13 16d = -80 d = -80/16 d = -5 a(35) = a(1) + (35 - 1)d = 13 + 34(-5) = 13 - 170 = -157

OpenStudy (anonymous):

thank you soooo much!!!!

OpenStudy (nurali):

Anytime.